# Robust Deterministic Policy Gradient in Zero-Sum Games for Disturbance Attenuation

### v5

@@ 4.2 Robust deep deterministic policy gradient
To implement RDPG in high-dimensional continuous control tasks, we incorporate the techniques from twin-delayed deep deterministic policy gradient (TD3)[5], resulting in the RDDPG. In particular, we introduce the following networks:
@@
    The online actor network $\pi_{\theta}$ for the user and the online actor network $\mu_{\phi}$ for the adversary
@@
    The corresponding target networks $\pi_{\theta^{\prime}}$, $\mu_{\phi^{\prime}}$ for the two online actor network
@@
    The corresponding target networks $Q_{\psi_{1}^{\prime}}(x,u,w)$, $Q_{\psi_{2}^{\prime}}(x,u,w)$
Now, following [5], the actor and critic networks are trained through the following procedures.
@@ 4.2.1 Critic update
$L_{\rm critic}(\psi_{i};B):=$ $\frac{1}{|B|}\sum\limits_{(x,u,w,c,x^{\prime})\in B}{(y-Q_{\psi_{i}}(x,u,w))}^{2},\quad i\in\{1,2\}$
where $B$ is the mini-batch, $|B|$ is the size of the mini-batch, $c$ is the cost incurred at the same time as the state $x$, and $x^{\prime}$ means the next state. Moreover, the target $y$ is defined as
@@
where $\epsilon_{1},\epsilon_{2}$ are random noise vectors added in order to smooth out the target values, $\mathcal{N}(0,\sigma)$ implies the Gaussian distribution with zero mean and variance $\sigma$, and the clip function is added in order to guarantee the boundedness of the control inputs. The critic’s online parameters $\psi_{1},\psi_{2}$ are updated by the gradient descent step to minimize the loss
@@
where $\alpha_{\rm critic}$ is the step-size. After the online critics are updated, the target parameters of two critic $\psi^{\prime}_{1}$ and $\psi^{\prime}_{2}$ are updated as follows:
${\psi^{\prime}_{i}}\leftarrow\tau{\psi_{i}}+(1-\tau){\psi^{\prime}_{i}},\quad i\in\{1,2\}$
where $\tau\in(0,1)$ serves as the interpolation coefficient, facilitating a gradual update that enhances training stability.
@@ 4.2.2 Actor update
Please note the opposite signs of the gradients for the user and adversary updates, which are due to their opposite roles. After the online actors are updated, the target parameters of actor $\theta^{\prime}$ and $\phi^{\prime}$ are updated as follows
$\theta^{\prime}\leftarrow\tau\theta+(1-\tau)\theta^{\prime},\quad\phi^{\prime}\leftarrow\tau\phi+(1-\tau)\phi^{\prime}$
where $\tau\in(0,1)$ is the interpolation coefficient.
4.2.3 Exploration
For exploration during training, Gaussian noise is added to both the user’s control action and the adversary’s disturbance. Specifically, the user’s action $u_{k}$ and the adversary’s disturbance $w_{k}$ at time step $k$ are expressed as
$u_{k}=$ $\pi_{\theta}(x_{k})+e_{1}$
$w_{k}=$ $\mu_{\phi}(x_{k})+e_{2}$
where $e_{1},e_{2}\sim\mathcal{N}(0,\sigma)$ represent zero-mean Gaussian exploration noises with standard deviation $\sigma$ [5]. The overall algorithm is summarized in Algorithm 1.
@@ 5 Quadrotor application
5.1 Quadrotor dynamics model
The quadrotor dynamics can be described as follows:
$m\mathbf{a}_{k}$ $=m\begin{bmatrix}0\\ 0\\ -g\end{bmatrix}+\mathbf{R}\mathbf{T}_{k}+\mathbf{F}_{D}+w_{k}$ (5)
$\Delta\omega_{k}$ $=I^{-1}(-\omega_{k-1}\times I\omega_{k-1}+\tau_{B})$ (6)
$\mathbf{v}_{k}$ $=\mathbf{v}_{k-1}+T_{s}\mathbf{a}_{k}$ (7)
$\mathbf{p}_{k}$ $=\mathbf{p}_{k-1}+T_{s}\mathbf{v}_{k}$ (8)
$\omega_{k}$ $=\omega_{k-1}+T_{s}\Delta\omega_{k}$ (9)
$\rho_{k}$ $=\rho_{k-1}+T_{s}\omega_{k}$ (10)
where $m$ is the mass of the quadrotor, $g$ is the gravity acceleration, $T_{s}$ is the sampling time, $w_{k}=[w_{k,x},w_{k,y},w_{k,z}]^{T}$ means an external disturbance applied to the quadrotor, $\mathbf{a}_{k}=[a_{k,x},a_{k,y},a_{k,z}]^{T}$ denotes linear acceleration, $\mathbf{v}_{k}=[v_{k,x},v_{k,y},v_{k,z}]^{T}$ represents linear velocity, $\mathbf{p}_{k}=[p_{k,x},p_{k,y},p_{k,z}]^{T}$ describes the position of the quadrotor in the inertial frame, $\Delta\omega_{k}=[\Delta\omega_{k,\phi},\Delta\omega_{k,\theta},\Delta\omega_{k,\psi}]^{T}$ is angular acceleration, $\omega_{k}=[\omega_{k,\phi},\omega_{k,\theta},\omega_{k,\psi}]^{T}$ is the angular velocity, and $\rho_{k}=[\phi_{k},\theta_{k},\psi_{k}]^{T}$ represents the roll, pitch, yaw of the quadrotor in the body frame. Moreover, $\mathbf{R}$ is the rotation matrix from the body frame to the inertial frame, and $\mathbf{T}_{B}$ is the total thrust of the quadrotor along the $z$ axis in the body frame and defined as follows:
$\mathbf{T}_{k}=\sum_{i=1}^{4}T_{k,i}=k_{F}\begin{bmatrix}0\\ 0\\ \sum\text{RPM}_{k,i}^{2}\\ \end{bmatrix}$
where $T_{k,i}$ is the thrust generated by $i$th motor, $k_{F}$ is the thrust coefficient, and $\text{RPM}_{k,i}$ is the speed of $i$th motor at time step $k$. $\mathbf{F}_{D}$ is a drag force, which is generated by motor spinning, defined as
$\mathbf{F}_{D}=-\mathbf{K}_{D}\left(\sum_{i=1}^{4}\text{RPM}_{k,i}\right)\mathbf{v}_{k}$
where $\mathbf{K}_{D}=\text{diag}(K_{D,x},K_{D,y},K_{D,z})$ is a matrix of drag coefficients. Among the four motors, two opposing motors ($1$th and 33rd) rotate clockwise, while the others (22nd and $4$th) rotate counterclockwise. This causes torque $\tau_{k}$ in the roll, pitch, yaw directions
$\tau_{k}=\begin{bmatrix}\tau_{k,\phi}\\ \tau_{k,\theta}\\ \tau_{k,\psi}\end{bmatrix}=\begin{bmatrix}l^{\prime}k_{F}(\text{RPM}_{k,2}^{2}-\text{RPM}_{k,4}^{2})\\ l^{\prime}k_{F}(-\text{RPM}_{k,1}^{2}+\text{RPM}_{k,3}^{2})\\ k_{M}(\text{RPM}_{k,1}^{2}-\text{RPM}_{k,2}^{2}+\text{RPM}_{k,3}^{2}-\text{RPM}_{k,4}^{2})\\ \end{bmatrix}$
where $l^{\prime}=\cos(\frac{\pi}{4})l$ is the scaled arm length, because the motors of the quadrotor are positioned at a 45-degree angle relative to the $x$ and $y$ axes when the yaw angle is 0, $k_{F}$ is the thrust coefficient, and $k_{M}$ is the torque coefficient. The inertia matrix, denoted by the symbol $I$, represents the inertial properties of the quadrotor’s body.
@@ 5.2 State space
$x_{k}=[\mathbf{e}_{k}^{p},\mathbf{e}_{k}^{v},\boldsymbol{\rho}_{k},\boldsymbol{\omega}_{k}]^{T}\in\mathbb{R}^{12}$
where $\mathbf{e}_{k}^{p}$ denotes the position error, $\mathbf{e}_{k}^{v}$ represents the velocity error, $\boldsymbol{\rho}_{k}$ is the Euler angle vector, and $\boldsymbol{\omega}_{k}$ is the angular velocity vector. These quantities are defined as follows:
@@
$=[p_{k,x}^{\text{target}}-p_{k,x},\ p_{k,y}^{\text{target}}-p_{k,y},\ p_{k,z}^{\text{target}}-p_{k,z}]^{T}\in\mathbb{R}^{3}$
@@
$=[v_{k,x}^{\text{target}}-v_{k,x},\ v_{k,y}^{\text{target}}-v_{k,y},\ v_{k,z}^{\text{target}}-v_{k,z}]^{T}\in\mathbb{R}^{3}$
$\boldsymbol{\rho}_{k}$ $=[\phi_{k},\theta_{k},\psi_{k}]^{T}\in\mathbb{R}^{3}$
$\boldsymbol{\omega}_{k}$ $=[\omega_{k,\phi},\omega_{k,\theta},\omega_{k,\psi}]^{T}\in\mathbb{R}^{3}$
where $\mathbf{p}_{k}^{\text{target}}$ and $\mathbf{p}_{k}$ denote the position vectors of the target and the quadrotor, respectively, and $\mathbf{v}_{k}^{\text{target}}$ and $\mathbf{v}_{k}$ represent their corresponding linear velocity vectors. $\phi_{k}$$\theta_{k}$$\psi_{k}$ denotes roll, pitch, yaw of the quadrotor. As in the method used in deep Q-network (DQN) [2], we used the last four time steps of the state as input to the neural network to help the network better capture dynamics in the environment.
@@ 5.3 Action space
The action of the quadrotor, $u_{k}$, is the speed of four motors $\text{RPM}_{k,i},\ i\in\{1,2,3,4\}$, which is continuous value between 0 and $21713.714$, as follows:
$u_{k}:=[\text{RPM}_{k,1},\text{RPM}_{k,2},\text{RPM}_{k,3},\text{RPM}_{k,4}]^{T}\in\mathbb{R}^{4}.$
The action of the disturbance adversary, $w_{k}$, represents the external forces to the quadrotor along the $x$, $y$, and $z$ axes of the body frame as follows:
$w_{k}:=[w_{k,x},w_{k,y},w_{k,z}]^{T}\in\mathbb{R}^{3}$
@@ 5.4 Cost function
The cost function $c(x_{k},u_{k},w_{k})$ is comprised of the cost function for the user $\tilde{c}(x_{k},u_{k})$ and the disturbance quadratic cost $\eta^{2}w_{k}^{T}w_{k}$. The cost function for the user $\tilde{c}(x_{k},u_{k})$ is the sum of the following terms: $\tilde{c}(x_{k},u_{k})=c_{k,p}+c_{k,v}+c_{k,\rho}+c_{l}+c_{tr}$
@@
    Position error cost $c_{k,p}$ : Position error between the target and quadrotor
@@
    Velocity error cost $c_{k,v}$ : Velocity error between the target and quadrotor
@@
    Angle cost $c_{k,\rho}$ : Magnitude of roll, pitch, and yaw of the quadrotor
@@
    Living cost $c_{l}$ : Constant cost at every step
    $c_{l}=\lambda$
  5. 5.
@@
where the scaling factors $\alpha$, $\beta$, $\epsilon$, and $\zeta$ are positive (i.e., $\alpha,\beta,\epsilon,\zeta>0$), and $\lambda$ is negative (i.e., $\lambda<0$). The specific values used for these parameters are summarized in the Appendix  2. Since $\alpha$ and $\beta$ are positive, the quadrotor agent learns a policy that minimizes position and velocity errors. If the quadrotor’s angles become too large, it becomes more susceptible to disturbances and may flip easily. Therefore, minimizing the angles cost $c_{\rho}$ is crucial for stability. Additionally, the living cost $c_{l}$ is a negative reward that encourages the quadrotor to maintain stability. The penalty cost $c_{tr}$ is a large positive value applied when the quadrotor fails to track the target—specifically, when the distance between the target and the quadrotor exceeds $5$ meters, or when the quadrotor flips with roll or pitch angles exceeding $\frac{\pi}{2}$. The agent learns a policy to avoid receiving this penalty, which can accelerate the initial learning process.
@@ 6.1 Experiments setups
For experiments, we used the gym-pybullet-drones environment [20], which is an open-source quadrotor simulator built with Python and the Bullet Physics engine. This environment provided a modular and precise physics implementation, supporting both low-level and high-level control. In the simulator, the Crazyflie 2.0 quadrotor was modeled, and its parameters were summarized in Appendix Table 1.
To demonstrate the robustness of the proposed RDDPG algorithm, we compared it with other model-free reinforcement learning algorithms, including DDPG [3], TD3 [5], and SAC [6]. These methods utilized a single agent that aimed to minimize the cost function $\tilde{c}$. To enhance their robustness against disturbances, random external disturbances were applied during the training process. A detailed summary of the hyperparameters used in training was provided in Appendix 3 and  4.
For evaluation, we generated two trajectories as shown in Fig. 2. During evaluation, we introduced random wind disturbances in the $x$ and $y$ directions. We considered three scenarios where the maximum wind velocity was limited to $\pm 5$, $\pm 10$, and $\pm 15$ [m/s], which were referred to as the $\pm 5$, $\pm 10$, and $\pm 15$ wind speed scenarios in the figures and tables. In each scenario, the initial wind speed was randomly set within a given range and updated every $10$ time steps with normally distributed noise.
Each algorithm was tested over 500 episodes and evaluated based on the mean and standard deviation of the total cost per episode, where the total cost was computed as the sum of the per-step costs $\tilde{c}(x_{k},u_{k})$ over a single episode. In addition, the mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor at each timestep were also considered as evaluation metrics.
@@ 6.2 Results
Figure 3: Mean and standard deviation of the total episode cost. RDDPG achieves the lowest average cost and the smallest variance across all scenarios. This indicates that the RDDPG yields a robust agent that performs consistently even under severe external disturbances.
[…]
@@
Figure 4: Trajectory results of all algorithms under the $\pm 15$ wind scenario and Traj (1).
Figure 4 illustrates the trajectory tracking performance of each algorithm across varying wind speeds. RDDPG achieves the most accurate trajectory tracking, exhibiting the smallest tracking error compared to the other algorithms. The mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor is presented in Figure 5. Across all scenarios, RDDPG consistently achieves lower position errors compared to TD3 and SAC. In the case of DDPG, it exhibits lower position errors under mild wind disturbance scenarios. However, as the disturbance becomes stronger, its performance degrades significantly, resulting in higher errors compared to RDDPG. Furthermore, DDPG shows larger standard deviations, indicating less stable performance across episodes. Detailed numerical results are provided in at Appendix Table 5.
Figure 5: Mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor at each time step. RDDPG consistently achieves low error and standard deviation across all scenarios, whereas other methods exhibit unstable performance.
@@ 7 Conclusion
This paper proposes RDPG, which combines the concept of the $H_{\infty}$ control problem and a two-player zero-sum dynamic game framework, to overcome the robustness problem of DRL algorithms. In RDPG, the user seeks to minimize the cost function while the adversary attempts to maximize it, thereby enabling the agent to learn a control policy that is robust to external disturbances. Furthermore, we introduce RDDOG, which integrates the robustness framework of RDPG with the learning stability of TD3, achieving both improved disturbance resistance and efficient policy learning in continuous control tasks. In order to evaluate the robustness of RDDPG, we implement it on quadrotor tracking tasks under various external disturbances. Experimental results show that RDDPG successfully learns an optimal control policy while maintaining stable performance across different disturbance scenarios. These findings demonstrate that RDDPG outperforms other DRL-based methods in terms of robustness and reliability in real-world-inspired environments.

### v6

@@ 4.2 Robust deep deterministic policy gradient
To implement RDPG in high-dimensional continuous control tasks, we incorporate the techniques from twin-delayed deep deterministic policy gradient (TD3) (TD3), resulting in the RDDPG. In particular, we introduce the following networks:
@@
    The online actor networks $\pi_{\theta}(x)$ and $\mu_{\phi}(x)$ for the user adversary respectively
@@
    The corresponding target actor networks $\pi_{\theta^{\prime}}(x)$, $\mu_{\phi^{\prime}}(x)$
@@
    The corresponding target critic networks
    $Q_{\psi_{1}^{\prime}}(x,u,w)$, $Q_{\psi_{2}^{\prime}}(x,u,w)$
Now, following TD3, the actor and critic networks are trained through the following procedures.
@@ 4.2.1 Critic update
$L_{\rm critic}$ $(\psi_{i};B):=$
$\frac{1}{|B|}\sum\limits_{(x,u,w,c,x^{\prime})\in B}{(y-Q_{\psi_{i}}(x,u,w))}^{2},i\in\{1,2\}$
where $B$ is the mini-batch, $|B|$ is the size of the mini-batch. Moreover, the target $y$ is defined as
@@
where $\epsilon_{1},\epsilon_{2}$ are random noise vectors added in order to smooth out the target values, $\mathcal{N}(0,\sigma)$ implies the Gaussian distribution with zero mean and variance $\sigma$, and the clip function is added in order to guarantee the boundedness of the control inputs with bound $\delta$. The critic’s online parameters $\psi_{1}$ and $\psi_{2}$ are updated by the gradient descent step to minimize the loss $L_{\rm critic}(\psi_{i};B)$
@@ 4.2.2 Actor update
Please note the opposite signs of the gradients for the user and adversary updates, which are due to their opposite objectives.
@@ 5.1 State space
$x_{k}=[\mathbf{e}_{k}^{p},\mathbf{e}_{k}^{v},\boldsymbol{q}_{k},\boldsymbol{\omega}_{k}]^{T}\in\mathbb{R}^{13}$
where $\mathbf{e}_{k}^{p}$ denotes the position error, $\mathbf{e}_{k}^{v}$ represents the velocity error, $\boldsymbol{q}_{k}$ is the quaternion vector, and $\boldsymbol{\omega}_{k}$ is the angular velocity vector. These quantities are defined as follows:
@@
$=[p_{x,k}^{\text{target}}-p_{x,k},\ p_{y,k}^{\text{target}}-p_{y,k},\ p_{z,k}^{\text{target}}-p_{z,k}]^{T}\in\mathbb{R}^{3}$
@@
$=[v_{x,k}^{\text{target}}-v_{x,k},\ v_{k,y}^{\text{target}}-v_{y,k},\ v_{z,k}^{\text{target}}-v_{z,k}]^{T}\in\mathbb{R}^{3}$
$\boldsymbol{q}_{k}$ $=[q_{w,k},q_{x,k},q_{y,k}q_{z,k}]^{T}\in\mathbb{R}^{4}$
$\boldsymbol{\omega}_{k}$ $=[\omega_{\phi,k},\omega_{\theta,k},\omega_{\psi,k}]^{T}\in\mathbb{R}^{3}$
where $\mathbf{p}_{k}^{\text{target}}$ and $\mathbf{p}_{k}$ denote the position vectors of the target and the quadrotor, respectively, and $\mathbf{v}_{k}^{\text{target}}$ and $\mathbf{v}_{k}$ represent their corresponding linear velocity vectors. As in the method used in deep Q-network (DQN) (DQN), we used the last four time steps of the state as input to the neural networks to help the network better capture dynamics in the environment.
@@ 5.2 Action space
The action of the quadrotor, $u_{k}$, is the speed of four motors $\text{RPM}_{i,k},\ i\in\{1,2,3,4\}$, which is continuous value between 0 and $21713.714$, as follows:
$u_{k}:=[\text{RPM}_{1,k},\text{RPM}_{2,k},\text{RPM}_{3,k},\text{RPM}_{4,k}]^{T}\in\mathbb{R}^{4}.$
The disturbance of the adversary, $w_{k}$, represents the external forces to the quadrotor along the $x$, $y$, and $z$ axes of the body frame of the quadrotor as follows:
$w_{k}:=[w_{x,k},w_{y,k},w_{z,k}]^{T}\in\mathbb{R}^{3}$
For training stability, each component of the disturbance was constrained to lie within the range [0.1,,0.1][-0.1,,0.1].
@@ 5.3 Cost function
The cost function $c(x_{k},u_{k},w_{k},x_{k+1})$ is comprised of the cost function for the user $\tilde{c}(x_{k},u_{k},w_{k},x_{k+1})$ and the disturbance quadratic cost $\eta^{2}w_{k}^{T}w_{k}$. The cost function for the user $\tilde{c}(x_{k},u_{k},w_{k},x_{k+1})$ is the sum of the following terms: $\tilde{c}(x_{k},u_{k},w_{k},x_{k+1})=c_{p,k}+c_{v,k}+c_{\rho,k}+c_{tr}$
@@
    Position error cost $c_{p,k}$ : Position error between the target and quadrotor
@@
    Velocity error cost $c_{v,k}$ : Velocity error between the target and quadrotor
@@
    Angle cost $c_{\rho,k}$ : Magnitude of roll $\phi_{k}$, pitch $\theta_{k}$, and yaw $\psi_{k}$ of the quadrotor
@@
where the scaling factors $\alpha$, $\beta$, $\epsilon$, and $\zeta$ are positive (i.e., $\alpha,\beta,\epsilon,\zeta>0$), and $\lambda$ is negative (i.e., $\lambda<0$). The specific values used for these parameters are summarized in the Appendix of arxiv version.
@@ 6.1 Experiments setups
Figure 2: Mean and standard deviation of the total episode cost. RDDPG achieves the lowest average cost and the smallest variance. This indicates that the RDDPG yields a robust agent that performs consistently even under external disturbances.
@@
For experiments, we used the gym-pybullet-drones environment (gym), which is an open-source quadrotor simulator built with Python and the Bullet Physics engine. This environment provided a modular and precise physics implementation, supporting both low-level and high-level control. In the simulator, the Crazyflie 2.0 quadrotor was modeled, and its parameters were summarized in the accompanying arXiv version of the paper.
To evaluate the robustness of the proposed RDDPG algorithm, we compare it against deterministic policy gradient–based reinforcement learning methods, including TD3 (TD3), which employs a single agent that seeks to minimize the cost function $\tilde{c}$. Beyond these standard single-agent baseline, we further benchmark RDDPG against robust reinforcement learning algorithms such as RARL (RARL), PR-MDP, and NR-MDP (AR-MDP). These robust methods are formulated under the TZDG framework, where the user and adversary aim to minimize and maximize the cost function, respectively. The full set of hyperparameters used in our experiments is provided in the accompanying arXiv version of the paper.
For evaluation, we generated two trajectories as shown in Fig. 3. To assess robustness to external perturbations, we conducted experiments in which the quadrotor was required to track these trajectories while being subjected to the disturbances $w_{x},w_{y},w_{z}$ were randomly sampled from the interval [0.2,0.2][-0.2,0.2]. Each algorithm was evaluated over 500 episodes and assessed using the mean and standard deviation of the total cost per episode, where the total cost was computed as the sum of the per-step costs $\tilde{c}(x_{k},u_{k},w_{k},x_{k+1})$ over a single episode.
@@ 6.2 Results
Figure 2 compares the performance of RDDPG with TD3, RARL (RARL), PR-MDP, and NR-MDP  (AR-MDP) across two evaluation trajectories under random disturbances. Across both trajectories experiments, RDDPG consistently demonstrates the most stable and reliable performance. Its cost distribution exhibits the smallest variance among all evaluated methods, indicating that the learned policy maintains robust tracking behavior even under random disturbances. In contrast, TD3 and RARL often achieve higher peak rewards but suffer from significantly larger variance, revealing high sensitivity to perturbations and unstable control behaviors. PR-MDP and NR-MDP show more moderate performance with reduced variability compared to TD3 and RARL, yet they still fall short of the consistency exhibited by RDDPG.
@@ 7 Conclusion
This paper proposes RDPG, which combines the concept of the $H_{\infty}$ control problem and a two-player zero-sum dynamic game framework, to overcome the robustness problem of DRL algorithms. In RDPG, the user seeks to minimize the cost function while the adversary attempts to maximize it, thereby enabling the agent to learn a control policy that is robust to external disturbances. Furthermore, we introduce RDDPG, which integrates the robustness framework of RDPG with the learning stability of TD3, achieving both improved disturbance resistance and efficient policy learning in continuous control tasks. In order to evaluate the robustness of RDDPG, we implement it on quadrotor tracking tasks under various external disturbances. Experimental results show that RDDPG successfully learns an optimal control policy while maintaining stable performance across different disturbance scenarios. These findings demonstrate that RDDPG outperforms other DRL-based methods in terms of robustness and reliability.